OpenStudy (anonymous):

How is this solved? (2a^2 - 4a + 2) / (3a^2 - 3)

OpenStudy (anonymous):

Can you explain more thoroughly? I'm confused on the steps.

OpenStudy (phi):

to simplify, you factor the top and bottom and divide out any common terms to factor the bottom, (3a^2 - 3) the first step is notice you can "factor out" a 3 from both terms. Can you do that ?

OpenStudy (anonymous):

Honestly, I'm not sure. I'm even more confused than I was before from the other answer...

OpenStudy (phi):

can you distribute the 3 in 3( a^2 -1) ? what do you get ?

OpenStudy (phi):

distribute the 3 means multiply each term inside the parens by 3

OpenStudy (anonymous):

3a^2 - 3 ?

OpenStudy (phi):

yes. notice that is what you have in your problem. factoring out the 3 changes the (3a^2-3) back to 3(a^2 -1) and "distributing the 3" changes 3(a^2-3) back to 3a^2-3 when you see 3a^2 - 3 you see there is a 3 in both terms, so you can "undistribute the 3" or factor it out anyways, we now have 3(a^2 -1) in the bottom

OpenStudy (phi):

now notice the (a^2 - 1) can be written as ( a^2 -1^2) because 1*1 is 1 why do we want to do that ? See http://www.regentsprep.org/Regents/math/ALGEBRA/AV6/Lfactps.htm

OpenStudy (phi):

a^2 -1 (or a^2 - 1^2 to show we have two "squares") is the "difference of two squares, and can be factored. Can you figure out the factors (from the web site)?

OpenStudy (phi):

match your problem to (a^2 -b^2) = (a+b)(a-b) you have (a^2 - 1^2) can you get the factors ?

OpenStudy (anonymous):

It would be 0? If I'm working this correctly.

OpenStudy (phi):

no we can't get a number because you have letters ... what you do is match (a^2 - 1^2) to (a^2 - b^2) so we can use the rule (a^2 -b^2) = (a+b)(a-b)